Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
berni - 22:51 Thursday 15 May 2025 (66797) Print this report
Operator Report - Afternoon shift

ITF found in Commissioning mode with Suzanne working on WE and NE mirror thermal camera imagines.

The planned activities could not be performed because the ITF could not be locked in a stable way, see: WE glitches. Valerio has been informed about this problem.

 

At around 18:30 UTC, in agreement with Michal, we decide to keep the ITF locked on the IR.

 

Suspension

The WE COIL H3 is close to saturation; Valerio has been informed.

 

SBE

SPRB vertical position recovered with step motors at 18:36 UTC.

 

Vacuum

The DMS was reporting a problem with WE cryotrap: I informed Antonio and he made a patch for it.

Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×